dukedadon:

An equation is shown below: 4(2x − 5) = 4 Part A: How many solutions does this equation have? (4 points) Part B: What are the solutions to this equation?

dukedadon:

@extrinix help meh plez ?

Extrinix:

So first off let's distribute it out \( 4(2x-5)=4\) \(8x-20=4\) Add 20 to both sides \(8x=24\) Divide by 8 on both sides \(x=3\) -- So there would be 1 solution. And that solution would be, \(x=3\).

dukedadon:

@extrinix wrote:
So first off let's distribute it out \( 4(2x-5)=4\) \(8x-20=4\) Add 20 to both sides \(8x=24\) Divide by 8 on both sides \(x=3\) -- So there would be 1 solution. And that solution would be, \(x=3\).
ty ty

Extrinix:

No problem! ^^